Your teacher is giving a test worth 250 points. There are 68 questions. Some questions are worth 5 points and the rest are worth 2 points. How many of each type of questions are on the test?
Answer by robertb(5830)

5x + 2y = 250
x+y = 68 ==> -2x - 2y = -136
Adding corresponding sides,
3x = 114 ==> x = 38 5-point questions, and
30 2-point questions.